OpenStudy (anonymous):

State the Fundamental Thereom of calculus and use it to integrate: e^(t^2) tan(sqrt(t) dt , from x to x^4. Please show work Kind regards...

OpenStudy (anonymous):

\[\int\limits_{x}^{x^4} e ^{t ^{2}} \tan \sqrt{t} \]

OpenStudy (anonymous):

using fundamental theorem of calculus \[\Large \frac{d}{dx}\int\limits_{v(x))}^{u(x)}f(t)dt=f(u(x))u'(x)-f(v(x))v'(x)\] here we have u(x)-x^4 v(x)=x can you do this ???

OpenStudy (anonymous):

@unseenoceans can you do this now?

OpenStudy (anonymous):

sorry sami, I have problems with my screen. it goes on and off. I know how to plug in the limits b and I am not sure after that.

OpenStudy (anonymous):

okay screen is back on ...lemme see.

OpenStudy (anonymous):

\[e ^{x ^{4^{2}}} \tan \sqrt{x ^{4}}.2x - e ^{x ^{2}}-\tan \sqrt{x}.1\]

OpenStudy (anonymous):

how does that look sami . does it need more simplification?

OpenStudy (anonymous):

erivative of u(x)=x^4 is u'(x)=4x^3 !!!!!!! you wrote 2x !

OpenStudy (anonymous):

derivative*

OpenStudy (anonymous):

okay...2x should be 4x^3 ...my bad

OpenStudy (anonymous):

yeah i just seen that...silly mistake...its 4am in Ireland. ;)

OpenStudy (anonymous):

also it is tan and not -tan

OpenStudy (anonymous):

hmmm at 4am :P there is another mistake find it !!

OpenStudy (anonymous):

That was a typo. I swear..

OpenStudy (anonymous):

ok ! i am too quite famous for typos here ask someone about me :P

OpenStudy (anonymous):

lol...so is that the solution then sami?..is it that easy....It looks monstrous but it really is not!

OpenStudy (anonymous):

just plug and chug and a bit of chain rule and that's it...Thanks for your help. You've been really great!

OpenStudy (anonymous):

you 're welcome:) btw if you want you can take\[\Large e^{x^2}\] common!

OpenStudy (anonymous):

yes and (e^x^4^2 whould read e^x^6 ....

OpenStudy (anonymous):

after common it should go down to e^x^4

OpenStudy (anonymous):

what abou tan(sqrt(x) can I take that too?

OpenStudy (anonymous):

it should get to \[\Large e^{x^6}\] because when you used x^4 at the time of substitution it should be \[\Large e^{t^2}=e^{(x^4)^2}\] \[\Large e^{x^8}\]

OpenStudy (anonymous):

yes, you're correct I used the wrong law \[a^{m} \times a^{n} = a ^{m+n} \]

OpenStudy (anonymous):

correct one is (a^m)n = a^mn

OpenStudy (anonymous):

yes !

OpenStudy (anonymous):

Anyway thanks you sami.....time for me to go bed now....I really appreciate your time and effort....good luck and see you around on openstudy my friend!

OpenStudy (anonymous):

you are welcome:)